There are 5.00x10^26 atoms/m3 in a container of argon gas (atomic mass number 40). The pressure in the container when the atoms have a rms speed of 350 m/s is:

A. 4.53x10^4 N/m2

B. 1.01x10^4 N/m2

C. 1.01x10^5 N/m2

D. 1.36x10^5 N/m2

E. None of the above.
Answer #1

ny p= 2 x Ž 3 Van X rus r Vannas u26 p = 1/3 s vam P = ģx m P = 7x seranie x(250) V 26 29 N/m P = 802xlo None of the above